OpenStudy (anonymous):

Use a sum or difference formula to find the exact value of the given expression. sin (− π/12)

OpenStudy (irishboy123):

\[\frac{1}{4} - \frac{1}{3} = \frac{3 - 4}{12} = -\frac{1}{12}\] that's a big clue